OpenStudy (anonymous):

How to approach this? sequence {X_n } is defined by x_1 = sqrt ( 2) and X_n + 1 = sqrt ( 2 + X_n ) . Show that X_n + 1 > X_n for all integers n> 1

OpenStudy (anonymous):

\[X _{n} \] is defined by \[X _{1}=\sqrt{2}\] and \[X _{n+1}=\sqrt{2+X _{n}}\] show that \[X _{n+1} > X_{n}\] for all n>1

OpenStudy (freckles):

so have you tried a proof by induction?

OpenStudy (anonymous):

How should I do that with this one?

OpenStudy (freckles):

First step is to show x(1)<x(2) Then assume x(n)<x(n+1) for some integer n last step is to show x(n+1)<x(n+2)

OpenStudy (anonymous):

Give some more hints please? Cause im just thinking recursive right now, and I dont know what xn is?

OpenStudy (freckles):

this is how your sequence x(n) is defined: \[x_n=\sqrt{2+x_{n-1}}\] well where x(1)=sqrt(2)

OpenStudy (freckles):

\[x_{n}>0 \text{ for integer } n>0 \\ \\ \text{ so showing } \\ x_n<x_{n+1} \text{ is equivalent to showing } x_n^2<x_{n+1}^2 \text{ assume for some integer } n \text{ we have } \\x_{n}^2<2+x_{n-1} <2+x_n=x_{n+1}^2 \\ \] \[\text{ so we want to show } x_{n+2}^2>x_{n+1}^2\]

OpenStudy (freckles):

hint write out what x^2(n+2) is

OpenStudy (freckles):

are you still there?

OpenStudy (anonymous):

yes, im just trying to understand writing it down

OpenStudy (freckles):

just in case you don't understand my non-latex I'm asking you to write \[x^2_{n+2} \text{ is } \\ \text{ and you really only know how to do in terms of the previous entry } \\ \text{ but still this is what I'm looking for }\]

OpenStudy (freckles):

\[x^2_{n+2}=2+x_{n+1}\] you do understand this is what I'm looking for?

OpenStudy (freckles):

and that we made the assumption that x(n+1)>xn earlier

OpenStudy (freckles):

\[x^2_{n+2}=2+x_{n+1}>2+x_n\] but guess what 2+x(n) is equal to?

OpenStudy (anonymous):

\[x^2_{n+2}=2+x_{n+2}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes plus one...

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

\[\sqrt{2}\]?

OpenStudy (anonymous):

Yea I cant see the pattern...

OpenStudy (freckles):

\[\text{ we assumed } x_n<x_{n+1} \\ \text{ we have } x^2_{n+2}=2+x_{n+1}>2+x_n=x_{n+1}^2\]

OpenStudy (freckles):

the end

OpenStudy (anonymous):

what do you mean with the end? do you have to the the induction steps with that?

OpenStudy (anonymous):

have to do*

OpenStudy (freckles):

I didn't give the whole proof. You still have to show x(2)>x(1). But yeah we did the whole assumption thing that x(n)<x(n+1) for some integer n then we showed x(n+2)>x(n+1) using our inductive assumption (hypothesis)

OpenStudy (freckles):

but the part you have to do isn't too hard it is just pluggin in n=1 into \[x_{n+1}=\sqrt{2+x_n}\]

OpenStudy (freckles):

then concluding x(2)>x(1) from that

OpenStudy (freckles):

I already did the other part of the proof

OpenStudy (freckles):

if you don't understand something I did let me know

OpenStudy (anonymous):

Yea I will try some on paper now, I will tag you here with questions :) thanks for now!

OpenStudy (anonymous):

Im putting in n=1 and gets that \[x _{2}=\sqrt{2+x _{1}}\] so \[x _{2}=\sqrt{2+\sqrt{2}}\] so \[x _{2}>x_{1}\] so we have prooven that the assumption is true for n=1 but the next step is to proove that its true for n=p and also true for p+1?

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

right which we have done above

OpenStudy (freckles):

we assumed \[x_n<x_{n+1} \text{ and it of course follows that } x_n^2<x_{n+1}^2 \\ \text{ and we need to show } x^2_{n+2}>x^2_{n+1}\] remember we did this: \[x_{n+2}^2=2+x_{n+1}>2+x_n=x_{n+1}^2\]

OpenStudy (freckles):

which you know from that we have x(n+2)>x(n+1) since x^2(n+2)>x^2(n+1)

OpenStudy (thomas5267):

As long as we prove that \(X_n>0\) we are done. Since \(a>b\implies\sqrt{a}>\sqrt{b}\), \(X_n>0\implies X_{n+1}=\sqrt{2+X_n}=\sqrt{2+\sqrt{2+X_{n-1}}}>\sqrt{2+X_{n-1}}={X_n}\)

OpenStudy (anonymous):

Yes but dont have to do it with all the right steps? 1) Inductionbase: you check if the assertion checks out for the lowest integer n. 2)Inductionassumption You assume that the assertion works with a integer n=p (Leftside_p=Rightside_p) 3)Inductionsteps: You show that if the assertion works with n=p it will work for n=p+1 also. I cant do that with this one...

OpenStudy (anonymous):

@thomas5267

OpenStudy (thomas5267):

Prove of \(X_{n+1}>X_n\): \[ X_1=\sqrt{2}\\ X_2=\sqrt{2+X_1}=\sqrt{2+\sqrt{2}}\approx \sqrt{3.41241}>X_1 \] Assume that \(X_n>X_{n-1}\), check \(X_{n+1}\) \[ \begin{align*} X_{n+1}&=\sqrt{2+X_n}\\ X_{n+1}^2&=2+X_n\\ X_n^2&=2+X_{n-1}\\ X_{n+1}^2-X_n^2&=2+X_n-2-X_{n-1}\\ &>2+X_n-2-X_n=0\quad\text{since }X_n>X_{n-1}\text{ by induction hypothesis}.\\ X_{n+1}^2&>X_n^2\\ X_{n+1}&>X_n \end{align*} \]

OpenStudy (anonymous):

what is happening here: \[X ^{2}_{n+1}-X ^{2}_{n}=2+X _{n}-2-X _{n-1}\] and then the next row\[>2+X _{n}-2-X _{n}=0\] where does the \[X _{n-1}\] go? @thomas5267

OpenStudy (thomas5267):

Replaced\(X_{n-1}\) with \(X_n\) and changed = to > since we assumed that \(X_n>X_{n-1}\). The induction hypothesis is the assumption.

OpenStudy (freckles):

just curious @pate16 what do you not understand about this: \[x_{n+2}^2=2+x_{n+1}>2+x_n=x_{n+1}^2\] given \[x_{n+1}>x_n\]

OpenStudy (freckles):

like you do know that: \[x^2_{n+2}=2+x_{n+1} \text{ and that } x_{n+1}^2=2+x_n \text{ right ? }\]

OpenStudy (freckles):

\[\text{ if } x_{n+1}>x_n \\ \text{ then } 2+x_{n+1}>2+x_n\]

OpenStudy (anonymous):

so this is what we have: \[X _{1}=\sqrt{2}\] \[X _{n+1}=\sqrt{2+Xn}\] and we want to show that \[X _{n+1}>X _{n}\] for alla integers n>1 1)(i.b) n=1 showed us that \[X _{2}>X _{1}\] 2)(steps) retriceuming that the assertion above is true for n=p \[X _{p+1}>X _{p}\] since \[X _{p+1}=\sqrt{2+X _{n}}\] it says \[\sqrt{2+X _{n}}>X _{p}\] we now want to show that the assertion applies to the nextcoming integer, i.e p+1 \[\sqrt{2+X _{p+1}}>Xp+1 \] substitution for x_p+1\[(\sqrt{2+\sqrt{2+X _{p}}})^{2}>(\sqrt{2+X _{p}})^2\] \[2+\sqrt{2+X _{p}}>2+X _{p}\] -2 both sides \[\sqrt{2+x _{p}}>x _{p}\] which shows that \[X _{n+1}>X _{n}\] cause its the same

OpenStudy (anonymous):

Is this correct? @freckles

OpenStudy (anonymous):

@IrishBoy123 you gotta help me out on this one! Im all stuck and it feels like the more I read the dumber I get

OpenStudy (irishboy123):

let me try paraphrase the advice @freckles gave do the plugging in bit of course at n = 1 or whatever , but at the heart of it is this idea assume \(x_{k+1} > x_k\) is true , ie for n = k we know from the definition that \(x_{k+2}^2 = 2 + x_{k+1}\) we also know that \(2 + x_{k+1} > 2 + x_{k}\) because we assumed \(x_{k+1} > x_k\) and we know that \(2 + x_{k} = x_{k+1}^2\) from the definition ***so we now know that assuming \(x_{k+1} > x_k\) is true , ie for n = k then \(x_{k+2}^2 >x_{k+1}^2 \) is also true

OpenStudy (irishboy123):

and if it's any consolation, these things give me a sore head too ;-)

OpenStudy (anonymous):

@IrishBoy123, so thats it with no calculation whatsoever? Is that just true..?

OpenStudy (freckles):

you do know in induction: we show the base case then we assume for some integer k blah blah is true then we show it is true for k+1 ?

OpenStudy (anonymous):

Yes

OpenStudy (freckles):

\[\text{ Base case } x_{2}=\sqrt{2+x_1}=\sqrt{2+\sqrt{2}}>\sqrt{2} \text{ since } 2+\sqrt{2}>2 \\\] \[\text{ Our assumption } x_{p+1}>x_p \text{ for some integer } p>1 \\ \text{( since you like to use the letter) } p\] \[\text{ What we want to show } x_{p+2}>x_{p+1}:\] \[\text{ Now remember our sequence is } x_{p+1}=\sqrt{2+x_p} \\ \text{ or squaring both sides } x^2_{p+1}=2+x_p \\ \text{ ok I'm going to replace } p \text{ with } p+1 \\ x^2_{p+1+1}=2+x_{p+1} \\ x^2_{p+2}=2+x_{p+1} \\ \text{ but remember in our assumption we got to assume } x_{p+1}>x_p \\ \text{ so we have } x^2_{p+2}=2+x_{p+1}>2+x_p \\ \text{ but remember that } x_{p+1}^2=2+x_p \text{ so we have } x^2_{p+2}=2+x_{p+1}>2+x_p=x^2_{p+1} \\ \text{ this says } x^2_{p+2}>x^2_{p+1} \\ \text{ which means } x_{p+2}>x_{p+1}\]

OpenStudy (freckles):

so it isn't just true we proved it is true by induction

OpenStudy (freckles):

so it is true I just didn't want you to think we weren't doing anything to prove it because we did

OpenStudy (irishboy123):

there's a really hacky way to do this because \(x_{n+1} = \sqrt{2 + x_{n}}\) we have \(2 = x_{n+1}^2 - x_n = x_{n+2}^2 - x_{n+1} \) \( x_{n+1} - x_n = x_{n+2}^2 - x_{n+1}^2 \) so if you assume the LHS is >0, so must the RHS etc

OpenStudy (anonymous):

Haha, I got it now!!! thanks alot everyone @freckles, @IrishBoy123 and @thomas5267 damn, I was just overthinking it I believe...

OpenStudy (irishboy123):

are you supposed to be learning induction or would any proof do?

OpenStudy (anonymous):

Im learning induction!!@IrishBoy123

OpenStudy (irishboy123):

well, i hope you enjoyed this induction into induction :-)

OpenStudy (anonymous):

haha yes, I dont know what I was thinking at first..im gonna close this one now, thanks again guys!